OpenStudy (anonymous):

Can anyone help me with 2 trig problem? Is a review question for a test :(

OpenStudy (anonymous):

Use the dot product to find the magnitude of u if u=(6,-5)

OpenStudy (dumbcow):

\[|u| = \sqrt{u*u} = \sqrt{6^2 + (-5)^2}\]

OpenStudy (anonymous):

i used the formula and i did not get an answer that match any answer choices.

OpenStudy (anonymous):

i dont know what i am doing wrong

OpenStudy (dumbcow):

36+25 = 61 magnitude is sqrt(61)

OpenStudy (anonymous):

oh i got it thank you so much, sorry for troubling you DX

OpenStudy (anonymous):

could you help me with another one? please

OpenStudy (dumbcow):

yw

OpenStudy (anonymous):

Find the magnitude and direction angle of v=-2i-6j. Round direction angle to nearest degree.

OpenStudy (anonymous):

for this question do you use the same formula like the last question?

OpenStudy (dumbcow):

yes find magnitude the same way for angle use \[\tan \theta = \frac{y}{x}\] where (x,y) is (-2,-6)

OpenStudy (anonymous):

So for the magnitude it would be sqrt 40?

OpenStudy (dumbcow):

correct, they may want it simplified though \[\sqrt{40} = 2 \sqrt{10}\]

OpenStudy (dumbcow):

also for the angle, make sure you are aware which quadrant "v" is in (-2,-6) is in 3rd quadrant so angle must be between 180 and 270

OpenStudy (anonymous):

so would the answer be ||v||= 4 sqrt10 ; theta = 244?

OpenStudy (dumbcow):

no angle is wrong

OpenStudy (anonymous):

uhm because for my answer choice only this choice have the correct magnitude

OpenStudy (anonymous):

choices*

OpenStudy (dumbcow):

use a calculator to find arctan(3) , then add 180 to put it in 3rd quadrant

OpenStudy (anonymous):

so theta will be 251

OpenStudy (dumbcow):

yes well if you round to nearest degree its 252

OpenStudy (anonymous):

oh ok, thank you so much for helping me out :D

OpenStudy (dumbcow):

yw:)